# Quantitative Comparison of Fine-Tuning Techniques for Pretrained Latent Diffusion Models in the Generation of Unseen SAR Image Concepts

### v1

@@ A.2 Training images generation
Figure 18: Study on the UNet, TE1 and TE2: Generated images (1024 per 1024 pixels) per category for 9 different models at epoch 8 (same seed training and evaluation)
@@
Figure 19: Refining with VAE vs <SAR> Embedding Learning: Generated images (1024 per 1024 pixels) per category for 9 different models (same seed training and evaluation)
In Figures 18 and 19, we generate images for each model with the same evaluation seed at epoch 8. Furthermore, Stable Diffusion XL behaves deterministically under fixed random seeds and identical training configurations, enabling controlled comparative experiments. We also use the same training seed to ensure that only the configuration changes across experiments. We use several prompts to generate our evaluation images as follows:
@@
A.3 Best model images generation
Figure 20: Generated images (1024 per 1024 pixels) from the heart-rose-2 model. Each image corresponds to a distinct textual prompt.
The images shown in Figure 20 were generated using our best-performing model, heart-rose-2. The corresponding prompts used for each image are listed below. (a) "A SAR image of a vast landscape with a rectangular structure in an airport surrounded by roads and patches of vegetation."

### v2

@@ Appendix B Training images generation
Figure B.1: Study on the UNet, TE1 and TE2: Generated images (1024x1024px at 40cm) per category for 9 different models at epoch 8 (same seed training and evaluation)
@@
Figure B.2: Refining with VAE vs <SAR> Embedding Learning: Generated images (1024x1024px at 40cm) per category for 9 different models (same seed training and evaluation)
In Figures B.1 and B.2, we generate images for each model with the same evaluation seed at epoch 8. Furthermore, Stable Diffusion XL behaves deterministically under fixed random seeds and identical training configurations, enabling controlled comparative experiments. We also use the same training seed to ensure that only the configuration changes across experiments. We use several prompts to generate our evaluation images as follows:
@@
Appendix C Best model images generation
The images shown in Figure 16 were generated using our best-performing model, heart-rose-2. The corresponding prompts used for each image are listed below.
(a) "A SAR image of a vast landscape with a rectangular structure in an airport surrounded by roads and patches of vegetation."